# The JADES Origins Field: A New JWST Deep Field in the JADES Second NIRCam Data Release

### v1

@@ 4.1 Science Motivation
While the investment of JWST imaging time in the JOF is tremendous, this field does not yet have the spectroscopic resources of the HUDF. The JOF lies in the flanking depth region of GOODS-S and CANDELS, rather than in the HUDF itself. GOODS-S has had substantial spectroscopic attention from ground-based observatories and the 3D-HST survey, but it does not have the ultra-deep data that the HUDF, particularly from the deep VLT/MUSE program. The JOF also has only limited overlap with the JWST FRESCO program. In December 2023, this area will be targeted with 4-5 NIRSpec MSA configurations from program 1286 and 1287, which will provide coverage of some high-priority objects.
The NIRCam grism has been shown to produce a compelling capability to measure strong [OIII] and H emission lines from galaxies at (Sun et al., 2023a; Kashino et al., 2023; Matthee et al., 2023a; Oesch et al., 2023; Wang et al., 2023c). Such observations would be an excellent pairing to the imaging, providing spectroscopic redshifts, line diagnostics, and dynamical measurements to this deep window on the universe. The NIRCam instrument team therefore has decided to devote 32 hours of Cycle 3 guaranteed time toward such a program (program ID 4540, PI: Eisenstein).
There are a number of key motivations from the science opportunity from these data. Some of these are generic to NIRCam grism data, but many are specific to the opportunities of this special field.
The core goal is to make make more complete use of the opportunity at that the [OIII] and H lines are often extremely strong. Observing only one filter leaves a redshift gap, but observing in both the F356W and F444W filters provides complete coverage at with these two strong lines. Further, at –6.6, the data provide access to both lines, which is useful for studying the metallicity of the galaxy populations and which nails the line identification.
At –6.6, these data will include both H and H, the ratio of which allows us to infer the amount of dust reddening, indeed free from the complicated path-loss corrections that result from the NIRSpec MSA.
At –5.2, we will measure H in the grism but will also measure [OIII] and H in the F250M, F300M, and F277W imaging. This will allow us to extend the measurement of the ratio of [OIII] to Balmer lines to a larger population of galaxies.
These same strong lines are usually detected in the medium-band imaging, where comparison between the filters allows one to measure the morphology in the line, as distinct from the rest-optical continuum. When combined with the grism data at a spectral resolution , one can break the degeneracy between the velocity field and the morphology, providing spatially resolved velocity measurements in some galaxies (Nelson et al., 2023).
Extending this opportunity, we note that a portion of these program 4540 observations overlap cycle 1 F444W grism observations from FRESCO (Oesch et al., 2023) but with a dispersion direction that differs by about 50. This will improve the velocity field measurements, as slitless spectroscopy always has a degeneracy when there is a velocity gradient across the object in the dispersion direction. It also provides another way to confirm the association of grism lines to the galaxies in the undispersed images.
The high spectral resolution of the NIRCam grism yields important dynamical information from the line widths of the emission lines. The widest lines can indicate galactic outflows or AGN broad lines, giving critical clues to galactic feedback and the role of black holes in galaxy growth (Matthee et al., 2023b; Maiolino et al., 2023a; Greene et al., 2023; Furtak et al., 2023; Harikane et al., 2023b; Kokorev et al., 2023).
The spectroscopy also creates important opportunities for the modeling of the detailed infrared SEDs. For , these spectroscopically detected galaxies will have their Balmer breaks in the progression of medium bands from program 3215; this will allow careful characterization of star formation histories. The detailed SED shapes around the Balmer and 4000Å breaks are key for detecting stars older than about 50 million years, but the effects are subtle and can be confused by degeneracies between the rest-frame colors and the photometric redshifts. Limits on the strong Balmer lines also allow one to clean the effects of the weaker Balmer lines around 4000Å from the continuum photometry. Finally, the strong lines themselves can be included in the galaxy spectral synthesis analyses, as a constraint on the most recent star formation.
Some of these galaxies will have very deep MIRI F770W data from program 1180, probing rest-frame 1–2 at these redshifts. This will provide more robust constraints on stellar populations and in particular on the stellar mass of these emission-line galaxies. We stress the opportunity to compare these stellar masses to the dynamical masses from resolved velocity profiles. Although the MIRI data will only detect a subset of the galaxies, that subset can be used to refine the analysis in the NIRCam bands to improve mass determinations overall.
Moving beyond the modeling of single galaxies, an exciting aspect of the NIRCam grism is the ability to detect a rich pattern of galaxy overdensities at these high redshifts. The FRESCO, EIGER, and ASPIRE data sets have shown beautiful results in this regard (Kashino et al., 2023; Witstok et al., 2023; Helton et al., 2023; Wang et al., 2023c; Herard-Demanche et al., 2023; Sun et al., 2023b). The patterns are much easier to find in three-dimensional mapping, and the ability of the grism to observe nearly every galaxy in an unbiased way has been crucial for getting to dense enough coverage to show these structures. In doing so, we can place the galaxies in their environmental context, connecting all of the imaging information about morphology, SEDs, and emission-line properties to environment. We judged this opportunity to measure this large-scale structure context so compelling that we opted to spread this grism data to cover the inter-module gap, even though programs 1210 and 3215 do not cover it, so that we can measure the environment of galaxies that fall near the edge of the imaging modules.
Indeed, Helton et al. (2023) presents the discovery of a substantial arcminute-scale galaxy overdensity with 43 emission-line galaxy members extracted from the south-west corner of the FRESCO data. Some of the galaxies fall exactly in the program 4540 footprint, which will extend the mapping beyond the FRESCO boundary. Our deeper data will better measure the dynamics of this forming structure through the [OIII] and H detections of these galaxies.
Finally, while the above focuses on the strong [OIII] and H lines, these grism data will net spectroscopic detections at lower redshift of the weaker rest-infrared lines, confirming redshifts to strong star-forming dwarfs and measuring lines whose equivalent widths are likely too small to claim from the imaging. We particularly note the ability to detect Paschen to (and Paschen to ), the comparison of which to the rest-UV continuum measured in the HST imaging in this field and H available in the F150W and F162M (and out to F250M, for P) filters can more directly reveal dust-obscured star formation. He I 10830Å and [S III] 9532Å lines can also be detected in vigorous cases, with line identification coming from the medium-band detections of the stronger lines at bluer wavelengths.
@@ 4.2 Short-wave Data
We use the opportunity of program 4540 to add deep F070W imaging to the JOF, the final NIRCam broad-band filter and 15th NIRCam filter overall. Given the mosaic pattern described below, this will reach exposure times of 16–68 ks, depending on overlaps, mildly shallower than but comparable to the 55 and 73 ks in F090W and F115W, respectively. This will extend the JWST coverage of the SEDs and produce optical imaging much deeper than the ACS imaging from GOODS and CANDELS.
These F070W data provide key enhancements for the core science of the program, as it will provide direct sensitivity to a fainter population of Ly break galaxies at –6. This will provide a UV-selected population to be compared to the emission-line selected population in [OIII] and H and to the characterization of Balmer/4000Å breaks from the medium-band data.
@@ 4.3 Implementation
Driven by the desire to cover most of the 3.1 to 5 m range over the full program 1210 footprint as well as to cover the intermodule gap and a small boundary for environment measurements, we design program 4540 for a tight 22 mosaic pattern as shown in Figure 5. The same data will give partial wavelength coverage in a mildly broader area; this bounding area is largely covered by JADES medium-depth NIRCam imaging.
@@
As redshifts in this field are so valuable, we aim to achieve a depth that is notably fainter than the FRESCO observations. In the overlap regions of the 4 tight pointings, we will reach a depth of over 9 hrs of integration time in each filter, compared to 2 hrs in FRESCO F444W. The F356W filter data is yet more sensitive because of lower zodiacal background. We expect to measure several hundred redshifts at , in addition to the lower redshift galaxies detected in lower luminosity redder lines.
@@
The grism data are undeflected at 3.95 , so the regions of complete F356W spectra and complete F444W spectra fall in different portions of the NIRCam field of view. With the column-direction grism (GRISMC), we use 2 rows of the mosaic stepped so that we can get complete coverage over the full JOF NIRCam imaging footprint. We then offset between F356W and F444W mosaics by so that we get complete 3.1 to 5 coverage over a consistent footprint. GRISMC is preferred to GRISMR for this purpose because the dispersion direction is the same in the two modules.
Following the pattern well-used in our earlier GTO program, we are using the DEEP8 readout mode to minimize data volume. We will use long integrations (6 or 7 groups, 1159–1375 sec) to minimize overheads and reduce the impact of detector noise. We use 6–8 dither positions in each of the mosaic positions to produce good pixel redundancy and achieve the desired exposure time The INTRAMODULEBOX pattern is tightest and most optimal for our purpose. Because of allocation constraints, we have slightly different exposure times per pointing, and we use longer exposures on the two F356W grism images with more coverage of the JOF.
@@
Although slitless spectroscopy often uses multiple dispersion directions to robustly connect lines to the zero-order location and hence get unique wavelengths, we do not do this here. With the sharp NIRCam imaging and accurate grism spectral tracing models, the associations of lines to galaxies detected in the imaging is often unconfused. More importantly, this field has exceptionally deep multi-band imaging, and these high-equivalent emission lines (essentially all of the cases) show up easily in the medium-bands from programs 1180, 1210, and 3215. This sharply reduces concerns about the ability to robustly match up the line detections to the undispersed images. Further, the NIRCam imaging yields very high quality photometric redshifts, which reduces the pressure to have a second line detection to produce an identification of the line and hence a redshift.
@@
Direct images are required by the slitless spectroscopy observing template. We use short exposures (5–6 groups of SHALLOW4, 257–311 sec) because this field already has far deeper data. We use the same LW filters, but change in SW to F150W and F200W to increase the signal-to-noise in the short exposures and to assure some infrared coverage in the outer portions of field where we may fall off of previous NIRCam imaging.
In detail, we found that layout of this program in a rigid mosaic was impossible given a hole in guide star coverage. Because of the two out-of-field direct images per pointing and the offset between the pointings of the two filters, there are a total of 24 separate visits. To achieve schedulability at a consistent position angle, we had to decouple the 8 grism positions and make small (10-20) level adjustments to find guide stars. Hence, the mosaic is an irregular grid.
@@ 4.4 Coordinated Parallels with MIRI
Program 4540 will also produce coordinated parallels for MIRI imaging in the F770W and F1500W filters. For each filter, these will reach a depth of about 2 hrs in each of 4 overlapping pointings, with a peak exposure time around 9 hrs. We choose to use F770W in the outermost 4 pointings, yielding a contiguous field of about by , while we use F1500W in the inner 4 pointings, yielding a field of by that is contained within the F770W mosaic. As with JADES, we use SLOWR1 readout mode to limit the data volume, and we use multiple integrations per NIRCam integration to mitigate saturation.
These F770W data will extend the study of galaxies and particularly their stellar mass utilizing rest-frame 1–2 measurements, while the the F1500W data will provide insight into the role of AGN in –5 galaxies, as the hot dust in the nuclei can outshine the Rayleigh-Jeans tail of the stellar population. Finding such sources may be the most powerful means to identify heavily obscured AGN at .
However, the location of this parallel field does fall into an area currently lacking in NIRCam and HST imaging. Of course, the extended CDF-S does have deep ground-based imaging (such as the deep 5-year stack from the Dark Energy Survey SNe program) as well as imaging from Spitzer, which can support near-term use. We expect that this field is likely to be imaged with NIRCam in the future, as its location is easy to reach as a parallel when NIRSpec is pointed at the exceptional JOF NIRCam imaging, indeed at the position angle such that the NIRSpec quadrants align to the NIRCam geometry.
@@ 5 JADES Second Imaging Data Release
In this paper, we present the second imaging data release of JADES, including the GOODS-S imaging observed in September and October 2022. The first release (Rieke et al., 2023) included the NIRCam imaging of the Deep Prime program on the HUDF. This second release includes the deep NIRCam parallel from program 1210 and the medium-depth mosaic that flanks these two imaging fields. Caveats regarding data quality and some missing data due to short circuits in the NIRSpec MSA and one skipped visit are described in Eisenstein et al. (2023).
Figure 6 shows a three-color image of the field. The total area of the JADES imaging (excluding the FRESCO-only region) is 67.7 arcmin, and the catalogs contain 94,000 distinct objects. A summary of the image reduction and catalog processing follows.
To create calibrated imaging and object photometry catalogs, the NIRCam data are reduced in the same manner as described in Rieke et al. (2023, hereafter R23) with a few modifications. For this release, version 1.11.4 of the STScI jwst pipeline is used but with a custom astrometric system as discussed in R23. The JWST calibration reference data file used during processing was jwst_1130.pmap, which adjusted the photometric zeropoints slightly relative to the first JADES data release. We have checked that the flux ratios of objects shared between the current data release and Rieke et al. (2023) are broadly consistent with the changes in photometric zeropoint.
@@
The exposure times achieved in this mosaic are shown in Figure 7. These show the deep parallel field of program 1210, along with the year 1 data in the Deep and Medium mosaics of program 1180. Year 2 observations will double the depth of the Deep mosaic, add flanking Medium depth around the Deep region, and extend the footprint with parallel NIRCam imaging from program 1286 and 1287 (Eisenstein et al., 2023). Figure 8 shows maps of the corresponding 5 sensitivity for F115W, F200W, and F356W in diameter apertures, with aperture corrections applied. These maps are computed by placing randomized apertures throughout the fully reduced mosaic images, after masking sources and surrounding pixels with significant flux. As reported by Rieke et al. (2023), the 1180 Deep region reaches F200W depth of mag AB.
@@
To enhance the multi-band space-based photometry available for sources, we also perform photometry on the Hubble Legacy Field HST/ACS F435W, F606W, F775W, and F814W mosaics and HST/WFC3 F105W, F125W, F140W, and F160W mosaics (Illingworth et al., 2016) after rectifying them to our astrometric frame. We reprocess and release JWST/NIRCam F182M and F210M mosaics from data acquired by the JWST Extragalactic Medium-band Survey (JEMS; Williams et al., 2023) and First Reionization Spectroscopically Complete Observations (FRESCO; Oesch et al., 2023) programs. We additionally reprocess and release JWST/NIRCam F430M, F460M, and F480M images from data acquired by JEMS and combine reprocessed JWST/NIRCam F444W exposures originally acquired by FRESCO with our F444W stack. We include photometry from all JWST filters in our released catalogs, where available.
@@
The construction of the photometric catalogs largely follows that described by Rieke et al. (2023). The LW channel images are stacked to create an inverse-variance-weighted signal-to-noise ratio image for performing detection. Catalogs are constructed from the resulting segmentations, with automated and manual masking of stars and diffraction spikes applied. SExtractor (Bertin & Arnouts, 1996) windowed positions are used as object centroids. Aperture photometry is measured using a range of circular apertures, two Kron apertures, and segmentations as discussed in Rieke et al. (2023). Sky background uncertainties are computed by placing random apertures throughout the images and directly from the jwst pipeline ERR image layers when available. Aperture corrections are computed from model point-spread-functions constructed following Ji et al. (2023). Flags are set to indicate bad pixels that may impact object segments, whether sources may be influenced by brighter neighbors, and whether stars identified in the image may impact them or if the objects themselves are consistent with stars.
@@
In addition to the photometry, we have computed photometric redshifts following the methods described in Rieke et al. (2023) and Hainline et al. (2023). In particular, these use the EAZY code (Brammer et al., 2008) with templates as described in Hainline et al. (2023) and incorporating percent-level adjustments to the photometric zeropoints based on the global median offset of each filter in a first-round of fitting. For convenience, the photometric redshift catalog is included as FITS extension of the photometry catalog.
@@
The images and catalogs are available on the Mikulski Archive for Space Telescopes (MAST) at https://archive.stsci.edu/hlsp/jades, with DOI 10.17909/z2gw-mk31. We will incorporate the 3215 imaging using this same pipeline in an incremental release.

### v2

@@ 4.1 Science Motivation
While the investment of JWST imaging time in the JOF is tremendous, this field does not yet have the deep spectroscopic attention that the HUDF has received.
The NIRCam instrument team therefore decided to devote 32 hours of Cycle 3 guaranteed time to a NIRCam slitless spectroscopy program (program ID 4540, PI: Eisenstein), augmenting the MSA coverage from programs 1286, 1287, and 3215.
The NIRCam grism has been shown to produce a compelling capability to measure strong [OIII] and H emission lines from galaxies at (F. Sun et al., 2023; D. Kashino et al., 2023; J. Matthee et al., 2023; P. A. Oesch et al., 2023; F. Wang et al., 2023). By observing in the F322W2, F356W, and F444W filters, program 4540 produced slitless spectroscopy from 2.4 to 5 m over most of the JOF, covering the [OIII] or H lines at . Of course, rest-infrared lines are also detected from brighter low-redshift galaxies as well. Early reductions indicate that these data have measured over 1400 redshifts across 0–9, nearly 100 per square arcminute.
Slitless spectroscopy is an exciting pairing to the JOF deep medium-band imaging, as it provides precise redshifts and line shape measurements to the lines inferred from the imaging. Precise redshifts in turn allow the medium-band photometry to make more accurate inferences about the continuum breaks, free from uncertainties as to where the weaker lines might fall. Indeed, the measurement of the line emission in the imaging breaks the degeneracy in slitless spectroscopy between astrometric position and wavelength, allowing the combination to measure velocity gradients within a galaxy. Further, line shapes are a powerful tool to identify Type I AGN and strong galactic outflows (J. Matthee et al., 2024; R. Maiolino et al., 2024; J. E. Greene et al., 2024; L. J. Furtak et al., 2024; Y. Harikane et al., 2023b; V. Kokorev et al., 2023).
A general census of strong line emitters also provides measures of galactic environment (D. Kashino et al., 2023; J. Witstok et al., 2024; J. M. Helton et al., 2024; F. Wang et al., 2023; T. Herard-Demanche et al., 2025; F. Sun et al., 2024) for studies of SED properties. Indeed, although this program was originally planned only to include F356W and F444W grism data, the presence of an obvious overdensity of F300M-excess sources in the JOF due to two overdensities at and 3.70 encouraged us to split the F356W time to include the F322W2 filter, so that we would have H coverage of these structures. In addition, J. M. Helton et al. (2024) presents the discovery of a substantial arcminute-scale galaxy overdensity with 43 emission-line galaxy members extracted from the south-west corner of the GOODS-S FRESCO data (P. A. Oesch et al., 2023). Combining with the Cycle-3 pure-parallel NIRCam grism survey SAPPHIRES (F. Sun et al., 2025; F. D’Eugenio et al., 2025b), Program 4540 has extended this mapping spatially and add [OIII] detections to the H coverage from FRESCO.
Further aspects of the data reduction of this grism data and resulting spectral catalog will be presented in Sun et al. (in prep.).
@@ 4.2 Short-wave Data
We use the opportunity of program 4540 to add deep F070W imaging to the JOF, the final NIRCam broad-band filter and 15th NIRCam filter overall. Given the mosaic pattern described below, this reaches exposure times of 16–68 ks, depending on overlaps, mildly shallower than but comparable to the 55 and 73 ks in F090W and F115W, respectively. This extends the JWST coverage of the SEDs and produces optical imaging much deeper than the ACS imaging from GOODS and CANDELS.
These F070W data provide key enhancements for the core science of the program, as it provides direct sensitivity to a fainter population of Ly break galaxies at –6. This yields a UV-selected population to be compared to the emission-line selected population in [OIII] and H and to the characterization of Balmer/4000Å breaks from the medium-band data.
@@ 4.3 Implementation
Driven by the desire to cover the 3 to 5 m range over the full program 1210 footprint as well as to cover the intermodule gap and a small boundary for environment measurements, we designed program 4540 for a tight 22 mosaic. As shown in Figure 5, the same data give partial wavelength coverage across 2.4–5.0 m in a mildly broader area; this bounding area is largely covered by JADES medium-depth NIRCam imaging.
@@
As redshifts in this field are so valuable, we aim to achieve a depth that is notably fainter than the FRESCO observations ( depth of unresolved emission line is erg s-1 cm-2 around 4.3 µm). In the overlap regions of the 4 tight pointings, we reach a maximum depth of hrs of integration time in the F444W filter, compared to 2 hrs in FRESCO F444W. We had originally planned for an equal amount of time in F356W, but later decided to split those exposures evenly between the F356W and F322W2 filters, noting that the F356W filter data is yet more sensitive because of lower zodiacal background. Preliminary analyses suggest that the data have reached a depth for unresolved emission lines of erg s-1 cm-2 around 4.3 µm (with the F444W filter at a median exposure time of 5.4 hrs); erg s-1 cm-2 around 3.6 µm (with the F356W filter alone at a median exposure time of 3.0 hrs); and erg s-1 cm-2 around 2.8 µm (with the F322W2 filter at a median exposure time of 3.0 hrs).
@@
The grism data are undeflected at 3.95 , so the regions of complete F356W spectra and complete F444W spectra fall in different portions of the NIRCam field of view. With the column-direction grism (GRISMC), we use 2 rows of the mosaic stepped so that we can get complete coverage over the full JOF NIRCam imaging footprint. One observation observed in all 3 filters; this produces coverage maps that are displaced as a function of wavelength. The other 2 observations were offset, with F444W toward the northwest and F356W/F322W2 toward the southeast, so that we get complete 3 to 5 coverage over the JOF. GRISMC is preferred to GRISMR for this purpose because the dispersion direction is the same in the two modules. Blueward of 3.1 microns, the F322W2 data (which use the same pointing as F356W) sample an even smaller field, with gaps opening in the footprint at the very blue end.
We use 6 dither pointings in each of the mosaic pointings, using the INTRAMODULEBOX pattern. Because of allocation constraints, we have different exposure times. Two of the pointings use DEEP8 readout, with 5 groups for F322W2 and F356W and 8 groups for F444W. The other two pointings use 8 groups of MEDIUM8 readout, with F444W being observed twice.
@@
Although slitless spectroscopy often uses multiple dispersion directions to robustly connect lines to the zero-order location and hence get unique wavelengths, we do not do this here. With the sharp NIRCam imaging and accurate grism spectral tracing models, the associations of lines to galaxies detected in the imaging is unconfused in most cases. More important, because of the deep multi-band imaging, the high-equivalent emission lines (essentially all of the cases) show up easily in the medium-bands from programs 1180, 1210, and 3215. This sharply reduces concerns about the ability to robustly match up the line detections to the undispersed images. Combining with the emission-line maps from imaging can even break the degeneracy between morphology and velocity fields inherent to slitless spectroscopy (e.g., E. Nelson et al., 2024; A. L. Danhaive et al., 2025). Further, the NIRCam imaging yields very high quality photometric redshifts, which reduces the pressure to have a second line detection to produce an identification of the line and hence a redshift. That said, the lack of multiple dispersion angles does sharply limit the ability to robustly extract continuum spectra from the grism data.
@@
Direct images are required by the slitless spectroscopy observing template. We use short exposures (6 groups of SHALLOW4, 311 sec) because this field already has far deeper data. We use the same LW filters, but change in SW to F150W and F200W to increase the signal-to-noise in the short exposures, to assure some infrared coverage in the southwest edge of field where we may fall in gaps of other NIRCam imaging, and to get a 3rd separate year of data for time-domain studies.
@@ 4.4 Coordinated Parallels with MIRI
Program 4540 also produced coordinated parallels for MIRI imaging in five filters. The longer grism observation with all 3 LW filters is used for F1500W and F2100W; this produces the tightest footprint about 2′ square with total depth of about 7 ks per filter . The other two observations use F770W and F1000W, producing a footprint centered on the first with total depth of about 5 ks per filter. We note that one of the four MIRI F1000W pointings was lost because of insufficient disk space on the spacecraft, resulting in reduced coverage and depth at the southeastern corner of the mosaics. All of the direct imaging use F560W, producing a mildly larger footprint; the increased sensitivity of F560W will compensate for the shallower depth of 1–2 ks. The footprints of coordinated MIRI parallels are also shown in Figure 5. The longer observation uses FASTR1 readout; the other two use SLOWR1 to limit the data volume. Aside from F560W, we use multiple integrations per NIRCam integration to mitigate saturation. In detail, the number of integrations and groups are 3 & 11, 3 & 11, 3 & 71, and 14 & 31 in F770W, F1000W, F1500W, and F2100W, respectively; F560W uses either 13 slow or 100 fast groups. The JWST ETC forecasts 5-sigma depths of 25.0-25.3, 25.5, 25.0, 24.5, and 22.9 AB mag in the F560W, F770W, F1000W, F1500W, and F2100W bands, respectively, assuming a point source with an 0.6′′ diameter aperture.
These F770W and F1000W data will extend the study of galaxies and particularly their stellar mass utilizing rest-frame 1–2 measurements, while the longer wavelength data will provide insight into the role of AGN in –5 galaxies, as the hot dust in the nuclei can outshine the Rayleigh-Jeans tail of the stellar population. Finding such sources may be the most powerful means to identify heavily obscured AGN at .
However, the location of this parallel field does fall into an area currently lacking in NIRCam and HST imaging. Of course, the extended CDF-S does have deep ground-based imaging (such as the deep 5-year stack from the Dark Energy Survey SNe program and the recently released Rubin DP2) as well as imaging from Spitzer, which can support near-term use. We expect that this field is likely to be imaged with NIRCam in the future, as its location is easy to reach as a parallel when NIRSpec is pointed at the exceptional JOF NIRCam imaging, indeed at the position angle such that the NIRSpec quadrants align to the NIRCam geometry.
@@ 5 JADES Second Imaging Data Release
In this paper, we present the second imaging data release of JADES, including the GOODS-S imaging observed in September and October 2022. The first release (M. J. Rieke et al., 2023) included the NIRCam imaging of the Deep Prime program on the HUDF. This second release includes the deep NIRCam parallel from program 1210 and the medium-depth mosaic that flanks these two imaging fields. Caveats regarding data quality and some missing data due to short circuits in the NIRSpec MSA and one skipped visit are described in D. J. Eisenstein et al. (2023).
Figure 6 shows a three-color image of the field. The total area of the JADES imaging (excluding the FRESCO-only region) is 67.7 arcmin2, and the catalogs contain 94,000 distinct objects. A summary of the image reduction and catalog processing follows.
To create calibrated imaging and object photometry catalogs, the NIRCam data are reduced in the same manner as described in M. J. Rieke et al. (2023, hereafter R23) with a few modifications. For this release, version 1.11.4 of the STScI jwst pipeline is used but with a custom astrometric system as discussed in R23. The JWST calibration reference data file used during processing was jwst_1130.pmap, which adjusted the photometric zeropoints slightly relative to the first JADES data release. We have checked that the flux ratios of objects shared between the current data release and M. J. Rieke et al. (2023) are broadly consistent with the changes in photometric zeropoint.
@@
The exposure times achieved in this mosaic are shown in Figure 7. These show the deep parallel field of program 1210, along with the year 1 data in the Deep and Medium mosaics of program 1180. Year 2 observations doubled the depth of the Deep mosaic, added flanking Medium depth around the Deep region, and extended the footprint with parallel NIRCam imaging from program 1286 and 1287 (D. J. Eisenstein et al., 2023). Figure 8 shows maps of the corresponding 5 sensitivity for F115W, F200W, and F356W in diameter apertures, with aperture corrections applied. These maps are computed by placing randomized apertures throughout the fully reduced mosaic images, after masking sources and surrounding pixels with significant flux. Such errors are preferred to formal errors derived from pixel-level uncertainties because of the correlations between pixels induced by the mosaicking. As reported by M. J. Rieke et al. (2023), the 1180 Deep region reaches F200W depth of mag AB.
@@
To enhance the multi-band space-based photometry available for sources, we also perform photometry on the Hubble Legacy Field HST/ACS F435W, F606W, F775W, and F814W mosaics and HST/WFC3 F105W, F125W, F140W, and F160W mosaics (G. Illingworth et al., 2016; K. E. Whitaker et al., 2019) after rectifying them to our astrometric frame. We reprocess and release JWST/NIRCam F182M and F210M mosaics from data acquired by the JWST Extragalactic Medium-band Survey (JEMS; C. C. Williams et al., 2023) and First Reionization Spectroscopically Complete Observations (FRESCO; P. A. Oesch et al., 2023) programs. We additionally reprocess and release JWST/NIRCam F430M, F460M, and F480M images from data acquired by JEMS and combine reprocessed JWST/NIRCam F444W exposures originally acquired by FRESCO with our F444W stack. We include photometry from all JWST filters in our released catalogs, where available.
@@
The construction of the photometric catalogs largely follows that described by M. J. Rieke et al. (2023). The LW channel images are stacked to create an inverse-variance-weighted signal-to-noise ratio image for performing detection. Catalogs are constructed from the resulting segmentations, with automated and manual masking of stars and diffraction spikes applied. SExtractor (E. Bertin & S. Arnouts, 1996) windowed positions are used as object centroids. Aperture photometry is measured using a range of circular apertures, two Kron apertures, and segmentations as discussed in M. J. Rieke et al. (2023). Sky background uncertainties are computed by placing random apertures throughout the images and directly from the jwst pipeline ERR image layers when available. Aperture corrections are computed from model point-spread-functions constructed following Z. Ji et al. (2024). Flags are set to indicate bad pixels that may impact object segments, whether sources may be influenced by brighter neighbors, and whether stars identified in the image may impact them or if the objects themselves are consistent with stars.
@@
In addition to the photometry, we have computed photometric redshifts following the methods described in M. J. Rieke et al. (2023) and K. N. Hainline et al. (2024a). In particular, these use the EAZY code (G. B. Brammer et al., 2008) with templates as described in K. N. Hainline et al. (2024a) and incorporating percent-level adjustments to the photometric zeropoints based on the global median offset of each filter in a first-round of fitting. For convenience, the photometric redshift catalog is included as FITS extension of the photometry catalog. A comparison to spectroscopic calibrators was plotted for DR1 in M. J. Rieke et al. (2023) and K. N. Hainline et al. (2024a), showing good performance in the aggregate. Of course, calibrations for faint populations are meager, and one should be careful about how outliers, particularly of unusual colors, might impact one’s results. We note that with the medium-band photometry, galaxies with strong rest-optical emission lines often have more robust photometric redshifts, so performance is not a simple function of continuum magnitude.
@@
The images and catalogs are available on the Mikulski Archive for Space Telescopes (MAST) at https://archive.stsci.edu/hlsp/jades, with DOI 10.17909/z2gw-mk31. We will incorporate the 3215 imaging using this same pipeline in an upcoming release.